A contractor agrees to build a dam 180ft long and 20fthigh12ftwide at the bottom and 8ftwide at the top for 9.75 a cubic yard Find his profit if his costs were 10000?
Answer by Alan3354(69443)

A contractor agrees to build a dam 180ft long and 20fthigh12ftwide at the bottom and 8ftwide at the top for 9.75 a cubic yard Find his profit if his costs were 10000?
------------------
Find the volume.
Vol = area of the trapezoid x the length
= L*h*(20+8)/2 = 180*20*14
= 50400 cubic ft
50400/27 = 1866.67 cubic yards
----
Price = Vol*$9.75 = $17864
- $10000 = $7864 profit